Find the value of k so that the given points are sqrt 82 units apart. (-5, -7), (k,2)
Please show work

Answer by stanbon(75887)

Find the value of k so that the given
points are sqrt 82 units apart. (-5, -7), (k,2)
-------
(k+5)^2 + (2+7)^2 = 82
--------------
k^2 + 10k + 25 + 81 = 82
----
k^2 + 10k + 25 = 1
(k+5)^2 = 1
k+5 = 1 or k+5 = -1
k = -4 or k = -6
=============
Cheers,
Stan H.